The current increases from 15 mA to 60 mA in a 50 mH inductor over 10 s. What is the induced emf across the inductor? (in V) Inductance is when a change in the current flowing through a circuit produces an induced emf voltage ε-L Δ1 /At that opposes the change L is inductance, I is current, t is time Submit Answer Incorrect. Tries 1/2 Previous Tries Post Discussion Send Feedback
Answer #1

V = L * dI / dt

= [ 50 * 10-3 * (60 - 15) * 10-3 ] / 10

emf induced across inductor = 2.25 * 10-4 V = 0.000225 V